# Computational bottlenecks for denoising diffusions

### v1

@@ 2. $(b)$
5 Proof of Theorem 1
5.1 Auxiliary lemmas
We will use the following lemmas, whose proofs are deferred to Appendix E.
Lemma 5.1.
@@ 5.1 Auxiliary lemmas
We state the following non-asymptotic result from [Pen12].
Lemma 5.2 (Theorem 3.1, [Pen12].).
@@
In Appendix E, we will use the last lemma to prove the bound below.
Lemma 5.3 (Alignment bound).
@@
We also use the following lemma, which is implied in the proof of Lemma 6.4.
Lemma 5.4.
@@ 5.2 Proof of Theorem 1
Using Lemma 5.1, we know that $|\langle\hat{\boldsymbol{v}}_{t},\boldsymbol{W}\hat{\boldsymbol{v}}_{t}\rangle|\leq 4\sqrt{k\log(n/k)}\cdot\sqrt{t}$ with probability at least $1-\binom{n}{k}^{-6}$, say. Further
@@
Therefore, by using Lemma 5.4 and $k\log(n/k)\gg n^{1/2}$, we get that with probability at least $1-C_{*}\exp(-n^{1/4})$, $|{\mathcal{A}}(\boldsymbol{w};C)|\leq\max\{\sqrt{k},k^{2}/n\}$ for some constant $C>0$, where
@@
By Lemma 5.3, we know that with probability at least $1-C_{*}\exp(-cn^{1/3})$ for some $C_{*},c>0$, $|\langle\boldsymbol{v}_{t_{0}(1+\delta)},\boldsymbol{u}\rangle|\geq c\delta^{2}$. Since $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|$ is stochastically increasing with $t$ (Fact F.2), we actually have that for any $t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}$, with the same probability $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq c\delta^{2}$.
@@
$\overline{\boldsymbol{m}}_{t_{1}}$ $:=\Delta\sum_{i=1}^{\ell_{1}}\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{(i-1)\Delta},(i-1)\Delta)\,,\,,\;\;\;\;\ell_{1}:=\lfloor t_{\mbox{\tiny\rm alg}}(1+\delta)/\Delta\rfloor\,.$ (29)
@@
$\tilde{\boldsymbol{y}}_{t}=\overline{\boldsymbol{m}}_{t_{1}}+\boldsymbol{B}_{t}\,,\;\;\;\;$ (30)
@@
$\|\overline{\boldsymbol{m}}\|_{F}\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}\,.$
@@
$\langle\hat{\boldsymbol{v}}_{t},\tilde{\boldsymbol{y}}_{t}\hat{\boldsymbol{v}}_{t}\rangle=\langle\hat{\boldsymbol{v}}_{t},\overline{\boldsymbol{m}}\hat{\boldsymbol{v}}_{t}\rangle+\langle\hat{\boldsymbol{v}}_{t},\boldsymbol{B}_{t}\hat{\boldsymbol{v}}_{t}\rangle\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}+\max_{\boldsymbol{v}\in\Omega_{n,k}}\langle\boldsymbol{v},\boldsymbol{B}_{t}\boldsymbol{v}\rangle\,.$ (31)
Using Lemma 5.1, we get that
@@
Using Lemma 6.3, we get that $\max_{t_{\ell}\leq t\leq t_{\ell+1}}\|\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}}\|_{{\mbox{\rm\tiny op}}}\leq 16\sqrt{(t_{\ell+1}-t_{\ell})n}=16$ with probability at least $1-2\exp(-32n)$. Taking a union bound over $\mathcal{T}_{n}$, we get that
@@
Following exactly the argument as in the proof of Theorem 2 (in particular, Subsection 6.3.2), we get that $\lambda_{1}(\boldsymbol{W}_{t})\leq t/3$ simultaneously for all $t\geq t_{2}$, with high probability. On this event, $\lambda_{1}(\boldsymbol{A}_{t})<t/2$ with high probability (because $t/6\gg(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}$). Hence, with high probability $\phi_{2}(\tilde{\boldsymbol{y}}_{t},t)=0$ and hence $\hat{\boldsymbol{y}}_{t}=\tilde{\boldsymbol{y}}_{t}$ for all $t\in[t_{2},\infty)$.
@@
6 Proof of Theorem 2
6.1 Properties of the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$
Proposition 6.1.
The proof of this proposition is standard, and will be presented in Appendix F. We note that the rate in Equation (34) gets slower the closer $k$ is to $n$; it is super-polynomial if $n\gg k\log n$.
Proposition 6.2.
6.2 Auxiliary lemmas
The following lemmas are needed for the analysis of the generated diffusion. Their proofs are deferred to Appendices G, H, I, J.
Lemma 6.3.
Let $\boldsymbol{W}_{t}$ be a ${\sf GOE}$ process. Then for each time $t_{0}\geq 0$,
Lemma 6.4.
Then for any $C>4$, we have
Lemma 6.5.
Lemma 6.6 (Concentration for deformed GOE model).
for all $n$ large enough.
Remark 6.1.
We only use Lemma 6.6 for the alignment $\langle\boldsymbol{v}_{1}(\boldsymbol{Y}),\boldsymbol{v}\rangle^{2}$.
6.3 Analysis of the diffusion process: Proof of Theorem 2
6.3.1 Analysis up to an intermediate time
Define $t_{\mbox{\tiny\rm between}}=n^{2}$. Following the same strategy with Section A.2, we will first show that $\hat{\boldsymbol{m}}=\boldsymbol{0}$ up to $t_{\mbox{\tiny\rm between}}$ with high probability by analyzing only the noise process (in short, if $\hat{\boldsymbol{m}}=\boldsymbol{0}$ always, our generated diffusion coincides with the noise process). Our strategy is of the same nature as that of Section A.2. Indeed, we will attempt to prove that $\hat{\boldsymbol{m}}=\boldsymbol{0}$ simultaneously for all $t$, with high probability. In this phase ($0\leq t\leq t_{\mbox{\tiny\rm between}}$), we will show that $|\hat{S}|<k/2$ (c.f. definition in Algorithm 1, lines 7, 8) for $0\leq t\leq t_{\mbox{\tiny\rm between}}$, with high probability ($\boldsymbol{v}_{t}$ is the top eigenvector of $\boldsymbol{A}_{t}$, c.f. Algorithm 1). Note that line 5 of Algorithm 1 is not relevant in this phase. We first show this for a sequence of time points $\{t_{\ell}\}_{\ell\geq 1}$, then control the in-between fluctuations. We can set $t_{1}$ to be any value in $[0,n/2)$, as the algorithm returns 00 if $t<t_{\mbox{\tiny\rm alg}}=n/2$ anyway. We denote the ${\sf GOE}$ process
We choose the following time points:
$t_{\ell}=\dfrac{n}{2}-1+\dfrac{\ell}{n^{4}}.$
To exceed $t_{\mbox{\tiny\rm between}}=n^{2}$, we will need $n^{6}$ values of $\ell$. By union bound from Lemma 6.4 (recall also the definition of the set $A(\boldsymbol{v};C)$ from this Lemma),
  • (i)
  • (ii)
Lemma 6.7 (Corollary 2.5, [NTV17]).
We then obtain that
with probability at least $1-p_{n}\geq 1-c_{0}n^{-8}$, from Lemma 6.7 and $\boldsymbol{W}_{t_{\ell}}\sim\sqrt{t_{\ell}}\boldsymbol{W}_{1}$. Now from Lemma 6.5, we know that with probability at least $1-2\exp(-32n)$,
$\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}_{t}-\lambda_{1}(\boldsymbol{W}_{t_{\ell}})\geq-16\sqrt{n(t_{\ell+1}-t_{\ell})}.$
$32\sqrt{\dfrac{t_{\ell+1}-t_{\ell}}{t_{\ell}}}\geq n^{-3/2}\sum_{i=8}^{n}w_{i}^{2}.$
A simple bit of algebra shows that
$A=\bigcup_{i=1}^{7}A(\boldsymbol{u}_{t_{\ell},i};C).$
From Lemma 6.4 and a union bound, that $|A|\leq 7\max\{\sqrt{k},k^{2}/n\}$ with probability at least $1-O(\exp(-(1/3)n^{1/4}))$. For every $j\in A^{c}$, we have
$|v_{tj}|\leq\sum_{i=1}^{n}|w_{i}|\cdot|u_{t_{\ell},i,j}|<\sum_{i=1}^{7}|w_{i}|\cdot\dfrac{C\sqrt{\log(n/k)}}{\sqrt{n}}+\sum_{i=8}^{n}|w_{i}|\leq\dfrac{C^{\prime}\sqrt{\log(n/k)}}{\sqrt{n}}+\dfrac{64}{\sqrt{n}}<C^{\prime\prime}\sqrt{\dfrac{\log(n/k)}{n}}$
$\varepsilon>\dfrac{C\sqrt{\log(n/k)}}{\sqrt{n/k}}$
6.3.2 Analysis to the infinite horizon
We will prove that simultaneously for all $t\geq n^{2}$, Algorithm 1 always terminates at line 5, or that $\lambda_{1}(\boldsymbol{W}_{t})\leq t/2$. Similar to Subsection 6.3.1, we choose the following sequence of time points for all $\ell\geq 1$:
$t_{\ell}^{(2)}=n^{2}+\ell-1$
Next we control the in-between fluctuations. From a simple modification of Lemma 6.3, we have
so that by union bound
Acknowledgements
Appendix A Proof of Theorem 3
A.1 Properties of the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$
Proposition A.1.

### v2

@@ 4 Reduction of estimation to diffusion-based sampling
5 All Lipschitz polytime algorithms fail
In Section 3 (Theorems 1, 2 and 3) we proved that there exist near-optimizers of the score matching objective that perform poorly. This means that, unless we optimize this objective to very high precision (and the sample size is large enough to approximate the population objective with similarly high precision), DS will fail.
However, we did not rule out the possibility that the optimal polytime drift $\hat{\boldsymbol{m}}$ will perform well.
We next show that this is not the case, under an additional assumption, namely that the drift $\hat{\boldsymbol{m}}(\,\cdot\,;t)$ is Lipschitz continuous for $t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}(n,k)$, where $t_{\mbox{\tiny\rm alg}}(n,k)$ is defined as in Eq. (9). We do so under a modification of the target distribution. Letting $\mu^{0}_{n,k}$ be the centered version of $\mu_{n,k}$ we define
$\overline{\mu}_{n,k}=\frac{1}{2}\,\delta_{\boldsymbol{0}}+\frac{1}{2}\,\mu^{0}_{n,k}\,.$ (16)
In words, with probability $1/2$ we let $\boldsymbol{x}=\boldsymbol{0}$ and with probability $1/2$ we draw $\boldsymbol{x}=\tilde{\boldsymbol{x}}-\mathbb{E}[\tilde{\boldsymbol{x}}]$, $\tilde{\boldsymbol{x}}\sim\mu_{n,k}$, a sparse rank-one matrix, as in previous sections.
Theorem 6.
Assume $\underline{k}_{n}$ exists as in Conjecture 3.2, and let $k=k_{n}$ be such that $\underline{k}_{n}\vee\sqrt{n}\leq k_{n}\ll n$ (moderately sparse regime). Let $\hat{\boldsymbol{m}}_{n}:{\mathbb{R}}^{n\times n}\times{\mathbb{R}}_{\geq 0}\to{\mathbb{R}}^{n\times n}$ be a polynomial time denoiser for such that the following holds for some $\delta=\delta_{n}\leq n^{-c_{*}}$ with $c_{*}>0$:
  1. 1.
    $\hat{\boldsymbol{m}}_{n}$ is nearly optimal, namely (for $\boldsymbol{y}_{t}=t\boldsymbol{x}+\boldsymbol{W}_{t}$, $\boldsymbol{W}_{t}$ standard Brownian motion)
    $\sup_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\Big{|}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}_{n}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}-\frac{1}{2}\Big{|}=o_{n}(n^{-1})\,,$ (17)
    $\sup_{t\leq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}_{n}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}=o_{n}(1)\,,$ (18)
    and further, for any $t\geq 0$ the MSE of $\hat{\boldsymbol{m}}_{n}$ smaller or equal than the MSE of $c(\lambda_{1}(\boldsymbol{y}_{t}))\hat{\boldsymbol{m}}_{n}(\boldsymbol{y}_{t},t)$ for any function $c(\,\cdot\,)$ of the maximum eigenvalue of $(\boldsymbol{y}_{t}+\boldsymbol{y}_{T}^{{\sf T}})/\sqrt{2}$, and than the MSE of $\boldsymbol{P}_{{\sf B}}\hat{\boldsymbol{m}}_{n}$, for $\boldsymbol{P}_{{\sf B}}$ the projection onto the unit ball.
  2. 2.
    $\hat{\boldsymbol{m}}_{n}(\,\cdot\,,t):{\mathbb{R}}^{n\times n}\to{\mathbb{R}}^{n\times n}$ is $C/t$-Lipschitz for some constant $C$ and all $t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}$.
Further assume that for any polynomial time computable test $\phi:{\mathbb{R}}^{n\times n}\to[0,1]$, $\sup_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\big{|}\mathbb{E}\{\phi(\boldsymbol{y}_{t})\}-\mathbb{E}\{\phi(\boldsymbol{W}_{t})\}|\leq o_{n}(1)$.
Then, there exists a constant $\gamma_{0}>0$ such that and any $\Delta>0$
$\inf_{t\in{\mathbb{N}}\cdot\Delta\cap[0,n^{1+\gamma_{0}}]}W_{1}(\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{t},t),\boldsymbol{x})\geq\frac{1}{2}-o_{n}(1)\,.$ (19)
We note that the error of polytime denoisers in Eq. (17) (and correspondingly, the sampling error of Eq. (19)) is $1/2$ instead of $1$ because the trivial denoiser (which always returns $\boldsymbol{0}$) achieves error $1/2$.
Theorem 6 does not rule out the possiblility that there exists a near-optimizer of score matching that violates the Lipschitz condition and samples well. However, in for $t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}$ accurate estimation is possible with Lipschitz algorithms, and indeed many natural methods follow in this class (e.g. neural nets with bounded number of layers and suitable operator norm bounds on the weigths.)
6 Numerical illustration
Figure 1: Generating sparse rank-one matrices $\boldsymbol{x}\sim\tilde{\mu}_{n,k}$ using denoising diffusions, for $n=350$, $k=20$. Left: MSE of various denoisers (vertical line corresponds to the algorithmic threshold $t_{\mbox{\tiny\rm alg}}$.) Right: Frobenius norms of generated samples.
The theory developed in the previous section yields a concrete prediction of the failure mode of DS when applied to the distribution $\tilde{\mu}_{n,k}=(1/2)\delta_{\boldsymbol{0}}+(1/2)\mu_{n,k}$ (with $\mu_{n,k}$ the law of $\boldsymbol{x}=\boldsymbol{u}\boldsymbol{u}^{{\sf T}}$, $\boldsymbol{u}\sim\operatorname{Unif}(B_{n,k})$). Namely (for large $n$, and $\sqrt{n}\ll k\ll n$):
1. Given sufficient model complexity and training samples, we expect the learnt denoiser $\hat{\boldsymbol{m}}_{n}(t,\,\cdot\,)$ to achieve MSE close to $1/2$ for $t<(1-\delta)t_{\mbox{\tiny\rm alg}}$, and close to 00 for $t>(1+\delta)t_{\mbox{\tiny\rm alg}}$.
2. We expect DS based on such a denoiser to generate samples concentrated around $\boldsymbol{0}$.
We tested these predictions in a numerical experiment. We considered three polytime denoisers:
  • $(a)$
    The spectral-plus-projection denoiser of Algorithm 1;
  • $(b)$
    A modification of the latter whereby the eigenvector calculation is replaced by $25$ iterations of power method;
  • $(c)$
    A learned graph neural network (GNN) (Scarselli et al., 2008; Kipf and Welling, 2016).
We carry out experiments with denoiser $(b)$ because $\ell$ iterations of power method can be approximated by an $\ell$-layers GNN. Hence, method $(b)$ provides a baseline for GNN denoisers.
Our GNN architecture uses node embeddings that are generated by $3$ iterations of the power method and $10$ message passing layers. Each message passing layer comprises of the ‘message’ and ‘node-update’ multi-layer perceptrons (MLPs), both of which are $2$-layer neural networks with LeakyReLU nonlinearity. We simply use the complete graph with self-loops for node embedding updates. We find that ‘seeding’ the node embeddings with iterations of power method is crucial for effective training.
During training of the denoiser, we sample time points $t$ as follows: choose a time threshold $t_{\star}$, and sample so that times $t>t_{\star}$ are picked with total probability $0.95$ (and times $t\leq t_{\star}$ are picked with total probability $0.05$). Within each interval $(0,t_{\star}]$ and $(t_{\star},T)$, times are chosen at random. This allows the neural network to initially prioritize learning in a low-noise regime. Several fine-tuning steps are taken, for which $t_{\star}$ is gradually decreased to refine the network on lower SNR.
Empirically, training directly with $10$ layers is difficult, due to its depth. We find that training initially with $7$ layers, then subsequently introducing the later layers results in more stable training.
We train such a network using $N=10000$ samples $\boldsymbol{x}_{i}$ from the distribution $\overline{\mu}_{n,k}$, and evaluate their MSE on $N_{\text{test}}=15000$ samples.
Figure 1, left frame, reports the MSE achieved by the three denoisers $(a)$, $(b)$, $(c)$ as a function of $t/n$, for $n=350$, $k=20$. As GNNs are permutation-equivariant, we are training on $\approx 1\%$ of all possible outcomes, for $n=350$ and $k=20$. We observe that the GNN denoiser outperforms both the spectral algorithm and its approximation via power iteration. However, none of the three approaches can overcome the barrier at $t_{\mbox{\tiny\rm alg}}=n/2$, while they perform reasonably well above that threshold. This confirms the prediction at point 1 above.
On the right, we plot the histogram of Frobenius norms of samples generated with the GNN denoiser. These values are close to 00, which confirms the prediction at point 2 above. By using $\|\,\cdot\,\|_{F}$ as a $1$-Lipschitz test function, we obtain that the Wasserstein distance between diffusion samples and the target distribution is at least $0.48$ (the asymptotic prediction from theory is $0.50$).
Acknowledgements
Appendix A Notations
We write $\boldsymbol{W}\sim{\sf GOE}(n)$ if $\boldsymbol{W}=\boldsymbol{W}^{{\sf T}}$ is a random symmetric matrix with $(W_{ij})_{i\leq j\leq n}$ independent entries $W_{ii}\sim{\mathsf{N}}(0,2)$, and $W_{ij}\sim{\mathsf{N}}(0,1)$ for $i<j$. We say that $(\boldsymbol{W}_{t}:t\geq 0)$is a ${\sf GOE}(n)$ process if $\boldsymbol{W}_{t}\in{\mathbb{R}}^{n\times n}$ is a symmetric matrix with entries above and on the diagonal $(W_{t}(i,j):i<j\leq n;W_{t}(i,i)/\sqrt{2}:i\leq n;t\geq 0)$ forming a collection of $n(n+1)/2$ independent Brownian motions.
We use $C,C_{i},c_{i},\dots$ to denote absolute constants, whose value can change from line to line.
Appendix B Proof of Theorem 2
B.1 Properties of the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$
Proposition B.1.
The proof of this proposition is standard, and will be presented in Appendix H. We note that the rate in Equation (21) gets slower the closer $k$ is to $n$; it is super-polynomial if $n\gg k\log n$.
Proposition B.2.
B.2 Auxiliary lemmas
The following lemmas are needed for the analysis of the generated diffusion. Their proofs are deferred to Appendices J, K, L, M.
Lemma B.3.
Let $\boldsymbol{W}_{t}$ be a ${\sf GOE}$ process. Then for each time $t_{0}\geq 0$,
Lemma B.4.
Then for any $C>4$, we have
Lemma B.5.
Lemma B.6 (Concentration for deformed GOE model).
for all $n$ large enough.
We only use Lemma B.6 for the alignment $\langle\boldsymbol{v}_{1}(\boldsymbol{Y}),\boldsymbol{v}\rangle^{2}$.
B.3 Analysis of the diffusion process: Proof of Theorem 2
B.3.1 Analysis up to an intermediate time
Define $t_{\mbox{\tiny\rm between}}=n^{2}$. Following the same strategy with Section D.2, we will first show that $\hat{\boldsymbol{m}}=\boldsymbol{0}$ up to $t_{\mbox{\tiny\rm between}}$ with high probability by analyzing only the noise process (in short, if $\hat{\boldsymbol{m}}=\boldsymbol{0}$ always, our generated diffusion coincides with the noise process). Our strategy is of the same nature as that of Section D.2. Indeed, we will attempt to prove that $\hat{\boldsymbol{m}}=\boldsymbol{0}$ simultaneously for all $t$, with high probability. In this phase ($0\leq t\leq t_{\mbox{\tiny\rm between}}$), we will show that $|\hat{S}|<k/2$ (c.f. definition in Algorithm 1, lines 7, 8) for $0\leq t\leq t_{\mbox{\tiny\rm between}}$, with high probability ($\boldsymbol{v}_{t}$ is the top eigenvector of $\boldsymbol{A}_{t}$, c.f. Algorithm 1). Note that line 5 of Algorithm 1 is not relevant in this phase. We first show this for a sequence of time points $\{t_{\ell}\}_{\ell\geq 1}$, then control the in-between fluctuations. We can set $t_{1}$ to be any value in $[0,n/2)$, as the algorithm returns 00 if $t<t_{\mbox{\tiny\rm alg}}=n/2$ anyway. We denote the ${\sf GOE}$ process
We choose the following time points:
$t_{\ell}=\dfrac{n}{2}-1+\dfrac{\ell}{n^{4}}.$
To exceed $t_{\mbox{\tiny\rm between}}=n^{2}$, we will need $n^{6}$ values of $\ell$. By union bound from Lemma B.4 (recall also the definition of the set $A(\boldsymbol{v};C)$ from this Lemma),
  • (i)
  • (ii)
  • (iii)
    Let $m$ be a (sufficiently large) constant integer with $1\leq m\leq n$. The first $m$ components of $\boldsymbol{w}$ take up $1-o(1)$ in $L_{2}$-norm by (i) and (ii) with overwhelming probability, from which we can simply use triangle inequality to upper bound $|A(\boldsymbol{v}_{t};C)|$ according to Lemma B.4 for $\boldsymbol{u}_{1},\cdots,\boldsymbol{u}_{m}$, which incurs only a constant factor of error probability, by union bound.
Lemma B.7 (Corollary 2.5, Nguyen et al. [2017]).
We then obtain that
with probability at least $1-p_{n}\geq 1-c_{0}n^{-8}$, from Lemma B.7 and $\boldsymbol{W}_{t_{\ell}}\sim\sqrt{t_{\ell}}\boldsymbol{W}_{1}$. Now from Lemma B.5, we know that with probability at least $1-2\exp(-32n)$,
$\boldsymbol{v}_{t}^{{\sf T}}\boldsymbol{W}_{t_{\ell}}\boldsymbol{v}_{t}-\lambda_{1}(\boldsymbol{W}_{t_{\ell}})\geq-32\sqrt{n(t_{\ell+1}-t_{\ell})}.$
$64\sqrt{\dfrac{t_{\ell+1}-t_{\ell}}{t_{\ell}}}\geq n^{-3/2}\sum_{i=8}^{n}w_{i}^{2}.$
A simple bit of algebra shows that
$A=\bigcup_{i=1}^{7}A(\boldsymbol{u}_{t_{\ell},i};C).$
From Lemma B.4 and a union bound, that $|A|\leq 7\max\{\sqrt{k},k^{2}/n\}$ with probability at least $1-O(\exp(-(1/3)n^{1/4}))$. For every $j\in A^{c}$, we have
$|v_{tj}|\leq\sum_{i=1}^{n}|w_{i}|\cdot|u_{t_{\ell},i,j}|<\sum_{i=1}^{7}|w_{i}|\cdot\dfrac{C\sqrt{\log(n/k)}}{\sqrt{n}}+\sum_{i=8}^{n}|w_{i}|\leq\dfrac{C^{\prime}\sqrt{\log(n/k)}}{\sqrt{n}}+\dfrac{128}{\sqrt{n}}<C^{\prime\prime}\sqrt{\dfrac{\log(n/k)}{n}}$
$\varepsilon>\dfrac{C\sqrt{\log(n/k)}}{\sqrt{n/k}}$
B.3.2 Analysis to the infinite horizon
We will prove that simultaneously for all $t\geq n^{2}$, Algorithm 1 always terminates at line 5, or that $\lambda_{1}(\boldsymbol{W}_{t})\leq t/2$. Similar to Subsection B.3.1, we choose the following sequence of time points for all $\ell\geq 1$:
$t_{\ell}^{(2)}=n^{2}+\ell-1$
Next we control the in-between fluctuations. From a simple modification of Lemma B.3, we have
so that by union bound
Appendix C Proof of Theorem 1
C.1 Auxiliary lemmas
We will use the following lemmas, whose proofs are deferred to Appendix G.
Lemma C.1.
@@ C.1 Auxiliary lemmas
We state the following non-asymptotic result from Peng [2012].
Lemma C.2 (Theorem 3.1, Peng [2012].).
@@
In Appendix G, we will use the last lemma to prove the bound below.
Lemma C.3 (Alignment bound).
@@
We also use the following lemma, which is implied in the proof of Lemma B.4.
Lemma C.4.
@@ C.2 Proof of Theorem 1
Using Lemma C.1, we know that $|\langle\hat{\boldsymbol{v}}_{t},\boldsymbol{W}\hat{\boldsymbol{v}}_{t}\rangle|\leq 4\sqrt{k\log(n/k)}\cdot\sqrt{t}$ with probability at least $1-\binom{n}{k}^{-6}$, say. Further
@@
Therefore, by using Lemma C.4 and $k\log(n/k)\gg n^{1/2}$, we get that with probability at least $1-C_{*}\exp(-n^{1/4})$, $|{\mathcal{A}}(\boldsymbol{w};C)|\leq\max\{\sqrt{k},k^{2}/n\}$ for some constant $C>0$, where
@@
By Lemma C.3, we know that with probability at least $1-C_{*}\exp(-cn^{1/3})$ for some $C_{*},c>0$, $|\langle\boldsymbol{v}_{t_{0}(1+\delta)},\boldsymbol{u}\rangle|\geq c\delta^{2}$. Since $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|$ is stochastically increasing with $t$ (Fact H.2), we actually have that for any $t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}$, with the same probability $|\langle\boldsymbol{v}_{t},\boldsymbol{u}\rangle|\geq c\delta^{2}$.
@@
$\bar{\boldsymbol{m}}_{t_{1}}$ $:=\Delta\sum_{i=1}^{\ell_{1}}\hat{\boldsymbol{m}}(\hat{\boldsymbol{y}}_{(i-1)\Delta},(i-1)\Delta)\,,\,,\;\;\;\;\ell_{1}:=\lfloor t_{\mbox{\tiny\rm alg}}(1+\delta)/\Delta\rfloor\,.$ (34)
@@
$\tilde{\boldsymbol{y}}_{t}=\bar{\boldsymbol{m}}_{t_{1}}+\boldsymbol{B}_{t}\,,\;\;\;\;$ (35)
@@
$\|\bar{\boldsymbol{m}}\|_{F}\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}\,.$
@@
$\langle\hat{\boldsymbol{v}}_{t},\tilde{\boldsymbol{y}}_{t}\hat{\boldsymbol{v}}_{t}\rangle=\langle\hat{\boldsymbol{v}}_{t},\bar{\boldsymbol{m}}\hat{\boldsymbol{v}}_{t}\rangle+\langle\hat{\boldsymbol{v}}_{t},\boldsymbol{B}_{t}\hat{\boldsymbol{v}}_{t}\rangle\leq(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}+\max_{\boldsymbol{v}\in\Omega_{n,k}}\langle\boldsymbol{v},\boldsymbol{B}_{t}\boldsymbol{v}\rangle\,.$ (36)
Using Lemma C.1, we get that
@@
Using Lemma B.3, we get that $\max_{t_{\ell}\leq t\leq t_{\ell+1}}\|\boldsymbol{W}_{t}-\boldsymbol{W}_{t_{\ell}}\|_{{\mbox{\rm\tiny op}}}\leq 16\sqrt{(t_{\ell+1}-t_{\ell})n}=16$ with probability at least $1-2\exp(-32n)$. Taking a union bound over $\mathcal{T}_{n}$, we get that
@@
Following exactly the argument as in the proof of Theorem 2 (in particular, Subsection B.3.2), we get that $\lambda_{1}(\boldsymbol{W}_{t})\leq t/3$ simultaneously for all $t\geq t_{2}$, with high probability. On this event, $\lambda_{1}(\boldsymbol{A}_{t})<t/2$ with high probability (because $t/6\gg(1-\gamma)t_{\mbox{\tiny\rm alg}}(1-\varepsilon)+(\gamma+\delta)t_{\mbox{\tiny\rm alg}}$). Hence, with high probability $\phi_{2}(\tilde{\boldsymbol{y}}_{t},t)=0$ and hence $\hat{\boldsymbol{y}}_{t}=\tilde{\boldsymbol{y}}_{t}$ for all $t\in[t_{2},\infty)$.
@@
Appendix D Proof of Theorem 3
D.1 Properties of the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$
Proposition D.1.